OpenStudy (anonymous):

Please Help!!! Will give medal,and follow slope-intercept form the population density of New Mexico increased at a relatively constant rate from 1980 to 1999. in 1985, the population density was about 11.62 people per square mile. in 1999, the population density was about 14.25 people per square mile. Write an equation that gives the population density (in square mile) as a function of the number of years since 1980. what was the population density in 1990. I already know the slope is 263/1400, I'm just confused on what to do next.

OpenStudy (tkhunny):

You have two points: (1985,11.62) and (1999,14.25) Let's see how you did on thehe slope: \(\dfrac{14.25-11.62}{1999-1985} = \dfrac{2.63}{14} = \dfrac{263}{1400}\) Very good. A line in slope intercept form looks like this: y = mx + b We know thee slope. y = (263/1400)x + b Now what?

OpenStudy (anonymous):

you would put one of the pairs in right?

OpenStudy (anonymous):

x and y

OpenStudy (anonymous):

So set up your ordered pairs: (1985,11.62) (1999, 14.25) Hint: Time is ALWAYS the domain number. Then do this:  Then pick a pair and the slope and plug them in and get b (y int). Then set up new equation so it has slope and y int in it.